Given the following information, what is the desired profit margin?

D/E = 2
Current profit margin = 10%
R = 0.65 Capital intensity ratio = 2
Desired sustainable growth rate = 15%
Answer #1

Desired sustainable growth rate = ROE * b

15% = ROE * 0.65

ROE = 15% / 0.65 = 23.08%

Capital Intensity Ratio = 1 / Asset Turnover Ratio

2 = 1 / Asset Turnover Ratio

Asset Turnover Ratio = 1 / 2 = 0.50

Financial Leverage = Total Assets / Total Equity

= [2 + 1] / 1 = 3

ROE = Profit Margin * Asset Turnover * Financial Leverage

23.08% = Profit Margin * 0.5 * 3

Profit Margin = 23.08% / 1.5 = 15.38%